# arXiv:2603.02404v1[math.OC] 02 Mar 2026

### v1

@@ IV Experiments
We clear capacity in two distinct auctions: one summer auction, from May 1 to October 30, and one winter auction, from November 1 to April 30. We solve the problems with hourly dispatch decisions on the corresponding time horizons of approximately $4300$ hours. To limit the size of the problem, we aggregate units of the same technology type when their capacity is below a prescribed threshold, taken as $100$ MW and $10$ MW. With a $100$ MW threshold, the system comprises $107$ units in total, including 10 energy storage units, $2$ solar units, $4$ wind units, and $91$ conventional units. With a 10 MW threshold, the system comprises $305$ units, including $17$ energy storage units, $10$ solar units, $9$ wind units, and $269$ conventional units. We use a Value of Lost Load of \$100,000100,000/MWh. According to market practice, we set a market cap based on the net Cost of New Entry, set to \$14.514.5/kW-month.
@@ V Results
Throughout the runs, we monitor the in-sample standard deviation of our $EUE$ estimator and stop when it is on the order of $25\%$ of the estimated $EUE$. On our system, this number is reached when approximately $20,000$ samples are used. We therefore use this total number of samples as a stopping criterion for our experiments. For the 100100MW aggregated system, this happens after 2 hours, and $5$ hours for the 1010MW aggregated system.
@@ V-A Convergence of the algorithm
Figure 1 reports convergence diagnostics for the SD algorithm at aggregation level $10$ and $100$ MW over $5$ and $2$ hours of wall-clock time, respectively. Panel 1a shows the model gap $\Delta_{k}$ as a function of the iteration index. For both aggregation levels, the gap decreases rapidly and falls below $10^{-4}$ after a few hundred iterations, indicating that the master problem is well solved relative to the current sample information. Panel 1b displays the incumbent objective value $\hat{f}_{k}(\hat{x}_{k})$; it decreases substantially during the first $\sim 100$ iterations, after which improvements are marginal and subsequent iterations produce small fluctuations around a stable level. The difference in objective value between aggregation levels is due to a different per-unit scaling. Panel 1c shows the corresponding estimated expected cost of unserved energy, which also stabilizes after the first $300$ iterations, and is the same for both aggregation levels. We observe that the lower-dimensional runs ($100$ MW aggregation) converge faster than the higher-dimensional runs.

### v2

# Stochastic Optimization for Resource Adequacy in Capacity Markets with Storage and Renewables

@@ IV Experiments
We clear capacity in two distinct auctions: one summer auction, from May 1 to October 30, and one winter auction, from November 1 to April 30. We solve the problems with hourly dispatch decisions on the corresponding time horizons of approximately $4300$ hours. To limit the size of the problem, we aggregate units of the same technology type when their capacity is below a prescribed threshold, taken as $100$ MW and $10$ MW. With a $100$ MW threshold, the system comprises $107$ units in total, including 10 energy storage units, 22 solar units, 44 wind units, and $91$ conventional units. With a 10 MW threshold, the system comprises $305$ units, including $17$ energy storage units, $10$ solar units, 99 wind units, and $269$ conventional units. We use a Value of Lost Load of \$100,000100,000/MWh. According to market practice, we set a market cap based on the net Cost of New Entry, set to \$14.514.5/kW-month.
@@ V Results
Throughout the runs, we monitor the in-sample standard deviation of our EUEEUE estimator and stop when it is on the order of $25\%$ of the estimated EUEEUE. On our system, this number is reached when approximately $20,000$ samples are used. We therefore use this total number of samples as a stopping criterion for our experiments. For the 100100MW aggregated system, this happens after 2 hours, and 55 hours for the 1010MW aggregated system.
@@ V-A Convergence of the algorithm
Figure 1 reports convergence diagnostics for the SD algorithm at aggregation level $10$ and $100$ MW over 55 and 22 hours of wall-clock time, respectively. Panel 1a shows the model gap $\Delta_{k}$ as a function of the iteration index. For both aggregation levels, the gap decreases rapidly and falls below $10^{-4}$ after a few hundred iterations, indicating that the master problem is well solved relative to the current sample information. Panel 1b displays the incumbent objective value $\hat{f}_{k}(\hat{x}_{k})$; it decreases substantially during the first $\sim 100$ iterations, after which improvements are marginal and subsequent iterations produce small fluctuations around a stable level. The difference in objective value between aggregation levels is due to a different per-unit scaling. Panel 1c shows the corresponding estimated expected cost of unserved energy, which also stabilizes after the first $300$ iterations, and is the same for both aggregation levels. We observe that the lower-dimensional runs ($100$ MW aggregation) converge faster than the higher-dimensional runs.